# Hidden Bose-Einstein Singularities in Correlated Electron Systems

### v1

@@ Abstract
Hidden singularities in correlated electron systems, which are caused by pair fluctuations of electron-electron or electron-hole bubbles obeying Bose-Einstein statistics, are unveiled theoretically. The correlation function of each pair fluctuation is shown to have a bound in the zero Matsubara frequency branch, similarly as the chemical potential of ideal Bose gases. Once the bound is reached, the self-energy starts to acquire a component proportional to Green’s function itself, i.e. the structure called one-particle reducible, to keep the correlation function within the bound. The singularities are closely related with, but distinct from, phase transitions with broken symmetries. Passing down through them necessarily accompanies a change in the single-particle density of states around the excitation threshold, such as the pseudo-gap behavior found here for the negative-$U$ Hubbard model above the superconducting transition temperature.
@@ I Introduction
Systems of identical particles with integer spins embrace a unique phase transition called Bose-Einstein condensation characterized by the emergence of macroscopic quantum coherence Cornell95 ; Ketterle97 ; Leggett06 , which may be regarded as realization of the singularity of the Bose distribution function. Specifically for noninteracting systems below the critical temperature, the chemical potential is pinned to the value of the lowest energy level with the singularity, on which particles accumulate macroscopically. The purposes of the present paper are to show that similar singularities are also present in the pair-fluctuation channels of fermion systems described in terms of the Bose distribution function, and to clarify some basic features of the hidden phases realized after the singularities have been reached.
The singularities are relevant to logarithmic contributions to the grand thermodynamic potential given diagrammatically in terms of closed electron-electron or electron-hole bubbles; see Fig. 1 below. The argument of the logarithm can be expressed in terms of boson Matsubara frequencies, and the vanishing of it in the zero-frequency branch has been identified as the second-order transition point in the perturbative treatments of using the bare Green’s function, known as the Thouless criterion for superconductivity Thouless60 ; NSR85 and Stoner criterion for ferromagnetism Stoner38 ; Blundell01 , for example; they are also identical with the mean-field estimations of the critical temperatures as well as those based on the random-phase approximations IKK63 ; BS66 ; DE66 . Making Green’s function self-consistent and incorporating terms other than the logarithmic contributions, however, each singular point generally becomes different from the corresponding transition point with a broken symmetry, as shown below both analytically and numerically. It also has not been clarified what happens after the singular points are reached and how to describe the phases theoretically. These are the issues to be studied here.
This paper is organized as follows. In Sect. 2, we elaborate the hidden singularities and introduce a procedure to handle them theoretically. In Sect. 3, we present numerical results on the negative-$U$ Hubbard model. In Sect. 4, we provide concluding remarks. We use the units of $\hbar=k_{\rm B}=1$.
@@ II.1 Luttinger-Ward functional
One of the theoretical frameworks suitable for the present purposes is the Luttinger-Ward expression of the grand thermodynamic potential $\Omega$ given as a functional of Green’s function $\hat{G}$ LW60 ; DDM64 ; DDM64-2 ; HRCZ07 ; Kita11 . The stationarity condition of $\Omega$ in terms of $\hat{G}$ yields a self-consistent equation for $\hat{G}$, which have been known by various names such as Dyson’s, the Schwinger-Dyson, or Dyson-Gor’kov equation. Being nonlinear in $\hat{G}$, it can describe ordered phases with spontaneous broken symmetries continuously from the normal state on the same footing. Indeed, the ordered phases are characterized by the emergence of some novel component in the matrix $\hat{G}$ with particle-hole and spin degrees of freedom, such as the anomalous Green’s function for superconductivity and spin polarization in $\hat{G}$ for ferromagnetism. This Luttinger-Ward functional with full particle-hole and spin degrees of freedom can be written as Kita11
@@
where $\beta$ is the inverse temperature, $\hat{G}_{0}$ is the noninteracting Green’s function, $\hat{\Sigma}$ is the self-energy, and Tr denotes trace over the arguments of $\hat{G}$. The stationarity condition of Eq. (1) with respect to $\hat{G}$ yields the Schwinger-Dyson equation
@@
These are formally the exact expressions, and practical calculations can be performed based on some approximate expression for $\Phi\!=\!\Phi[\hat{G}]$, obeying Eq. (2) just like the exact theory. The scheme of Eq. (2) has also been known as conserving approximations of Kadanoff and Baym KB62 or $\Phi$-derivable approximations Baym62 . The potential $\Omega$ as a functional of $\hat{G}$ is alternatively called effective action in relativistic quantum field theory CJT74 ; Weinberg95 .
@@ II.2 Normal state
To elucidate the hidden singularities that lies near phase transitions with broken symmetries, we specifically consider the negative-$U$ Hubbard model at a low electron density $\bar{n}\!=\!N/V$, which has negligible lattice anisotropy and definitely exhibit $s$-wave superconductivity at low temperatures. We study this system based on the fluctuation-exchange approximation for superconductivity (FLEX-S approximation), whose $\Phi$ is given diagrammatically as Fig. 1 Kita11 . We first focus on the normal state, where the series reduces to the $\Phi$ functional of the standard fluctuation-exchange (FLEX) approximation BSW89 ; BS89 ; Kita11 ,
@@
Here, $\vec{q}\equiv({\bf q},i\omega_{\ell})$ is the four momentum consisting of wavevector ${\bf q}$ and boson Matsubara frequency $\omega_{\ell}$, and functions $\chi_{GG}(\vec{q}\,)$ and $\chi_{G\bar{G}}(\vec{q}\,)$ are defined in terms of the normal-state Green’s function $G(\vec{k})$ and its time-reversed partner $\bar{G}(\vec{k})\!\equiv\!G(-\vec{k})$ generally by
@@
where $\vec{k}\!\equiv\!({\bf k},i\varepsilon_{n})$ with $\varepsilon_{n}$ denoting the fermion Matsubara frequency; this function satisfies
@@
as can be shown by a change of summation variable $\vec{k}$. The first term in Eq. (3) is the Hartree-Fock contribution, the second one is the particle-particle ladder series relevant to the superconducting fluctuations including the second-order contribution, and the third and fourth terms represent the spin and charge fluctuations, respectively, composed of the particle-hole ladder and ring diagrams BSW89 . The corresponding self-energy is obtained by the functional differentiation of Eq. (2); for the normal state that is symmetric in the spin and particle-hole spaces, the differentiation simplifies to $\Sigma_{\rm n}(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi_{n}}{\delta G(\vec{k})}$ in terms of Eq. (3), yielding BSW89 ; BS89 ; Kita11 ; Yanase03 ; Kontani13
@@
Now, we are ready to discuss the hidden singularities in full detail. With replacing $G$ by $G_{0}$ in Eq. (5), the conditions of the vanishing denominators at $\vec{q}=0$ in the second and third terms reproduce the Thouless criterion $U\chi_{G_{0}\bar{G}_{0}}(\vec{0})\!=\!1$ for superconductivity NSR85 and Stoner criterion $U\chi_{G_{0}G_{0}}(\vec{0})\!=\!1$ for ferromagnetism Blundell01 , respectively. However, the renormalized conditions $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ do not specify the phase transition points of respective broken symmetries. Indeed, by linearizing the self-consistent equation for the anomalous Green’s function $F$ in the FLEX-S approximation, we obtain the equation for the superconducting transition temperature $T_{\rm c}$ that the eigenvalue problem
@@
have a non-trivial anomalous self-energy $\Delta(\vec{k})$ due to the emergence of a zero eigenvalue, where $U^{\rm s}_{\vec{k}\vec{k}^{\prime}}$ is the pairing interaction with correlations given explicitly by Eq. (91) of Ref. Kita11, , i.e.,
@@
This condition for $T_{\rm c}$, which is also identical with the one that the superconducting correlation function diverges, is manifestly different from $U\chi_{G\bar{G}}(\vec{0})\!=\!1$. The same statement holds true for the case of ferromagnetism with $U\!>\!0$, as can be checked easily based on the consideration of incorporating the spin polarization in Green’s function. Hence, we should allow for the possibility that $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ are reached independently of the phase transitions with broken symmetries. The fact that $U\chi_{G_{0}\bar{G}_{0}}(\vec{0})\!=\!1$ and $U\chi_{G_{0}G_{0}}(\vec{0})\!=\!1$ specify the mean-field $T_{\rm c}$ suggests that the renormalized ones will be realized near the phase transition points with broken symmetries.
@@
$\frac{1}{\beta}\sum_{\bf q}\left\{\ln\left[1-U\chi_{G\bar{G}}(q,0)\right]+U\chi_{G\bar{G}}(q,0)\right\},$ (8)
where $U\chi_{G\bar{G}}(q,0)$ is real, as can be shown based on Eq. (4). Since Eq. (8) is part of the grand thermodynamic potential which is real, the inequality $11-U\chi_{G\bar{G}}(q,0)\!\geq\!$ 1 should hold. Put it another way, the series expansion of Eq. (8) does not converge for $U\chi_{G\bar{G}}(q,0)>1$. Thus, we are led to the conclusion that $U\chi_{G\bar{G}}(\vec{0})$ and $U\chi_{GG}(\vec{0})$ must have a common bound 1, at which they should be pinned after reaching it; we call $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ hidden singularities.
The singularities can be defined more generally beyond the FLEX approximation in terms of the logarithmic series of Fig. 1 with the bare interaction $U$ replaced by some renormalized one $\underline{\Gamma}$: the matrix argument of the logarithm $\ln(\underline{1}+\underline{\Gamma}\,\underline{GG})$ comes to have a zero eigenvalue in the zero Matsubara frequency branch.Kita11 ; DDM64-2 Thus, $U$ in our analysis may be regarded as a renormalized effective interaction.
@@
We propose a resolution to this issue within the variational framework of Eqs. (1) and Eq. (2) by making use of the method of the Lagrange multiplier so as not to affect $\Omega$ explicitly. Specifically for the negative-$U$ case where $U\chi_{G\bar{G}}(\vec{0})$ is relevant, we add to the $\Phi$ functional the following term that vanishes for $U\chi_{G\bar{G}}(\vec{0})=1$,
$\varDelta\Phi_{\lambda}^{\rm n}\equiv\lambda_{\rm n}N\left[1-U\chi_{G\bar{G}}(\vec{0})\right],$ (9)
@@
where the particle number $N$ is incorporated in the prefactor to make the Lagrange multiplier $\lambda_{\rm n}$ intensive; note that $\lambda_{\rm n}$ has the dimension of energy. Performing the differentiation of Eq. (2) with Eq. (9) added to $\Phi$, we obtain the expression of the self-energy for the pinned phase of $U\chi_{G\bar{G}}(\vec{0})=1$ in terms of Eq. (5) as
$\Sigma_{\lambda}^{\rm n}(\vec{k})=\Sigma_{\rm n}(\vec{k})+\lambda_{\rm n}N\frac{U}{V}G(-\vec{k}).$ (10)
@@
The additional second term stemming from Eq. (9) is characteristic of the pinned phase, which corresponds to the condensate fraction of the ideal Bose-Einstein condensation Leggett06 . It has the unusual structure called one-particle reducible (1PR), i.e. the structure that becomes disconnected upon cutting a line, which originates from the two-particle reducible structure of Eq. (9). The parameter $\lambda_{\rm n}$ in Eq. (9) enables us to control the value of $U\chi_{G\bar{G}}(\vec{0})$ and should be chosen in such a way as to meet $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ by solving Dyson’s equation
$G(\vec{k})\!=\!\bigl{[}G_{0}^{-1}(\vec{k})\!-\!\Sigma_{\lambda}^{\rm n}(\vec{k})\bigr{]}^{-1}$ (11)
self-consistently. This completes the formulation for the normal state.
@@
The parameter $\lambda_{\rm n}$ is finite only for the pinned phase of $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and expected to grow almost linearly as a function of temperature initially. This linear change in the structure of the self-energy need not be considered in performing the first derivatives of $\Omega$ because of the stationarity condition (2). However, it does bring singularities in the second derivatives of $\Omega$ such as the specific heat, similarly as those caused by the emergences of $\Delta(\vec{k})$ for superconductivity and the spin-polarized self-energy for ferromagnetism near $T_{\rm c}$. Thus, the point $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ may be regarded as specifying a kind of transition without apparent broken symmetries. Detailed studies on the observable singularities remain to be performed in the future.

### v2

@@ Abstract
Hidden singularities in correlated electron systems, which are caused by pair fluctuations of electron-electron or electron-hole bubbles obeying Bose-Einstein statistics, are clarified theoretically. The correlation function of each pair fluctuation is shown to have a bound in the zero Matsubara frequency branch, similarly to the chemical potential of ideal Bose gases. Once the bound is reached, the self-energy starts to acquire a component proportional to Green’s function itself, i.e., the structure called one-particle reducible, to keep the correlation function within the bound. The singularities are closely related to, but distinct from, phase transitions with broken symmetries. Passing down through them is necessarily accompanied by a change in the single-particle density of states around the excitation threshold, such as the pseudogap behavior found here for the negative-$U$ Hubbard model above the superconducting transition temperature.
@@ I Introduction
Systems of identical particles with integer spins undergo a unique phase transition called Bose-Einstein condensation characterized by the emergence of macroscopic quantum coherence Cornell95 ; Ketterle97 ; Leggett06 , which may be regarded as the realization of the singularity of the Bose distribution function. Specifically for noninteracting systems below the critical temperature, the chemical potential is pinned to the value of the lowest energy level with a singularity, at which particles accumulate macroscopically. The purposes of the present paper are to show that similar singularities are also present in the pair-fluctuation channels of fermion systems described in terms of the Bose distribution function and to clarify some basic features of the hidden phases realized after the singularities have been reached.
The singularities are relevant to logarithmic contributions to the grand thermodynamic potential given diagrammatically in terms of closed electron-electron or electron-hole bubbles (see Fig. 1 below). The argument of the logarithm can be expressed in terms of boson Matsubara frequencies, and the vanishing of it in the zero-frequency branch has been identified as the second-order transition point in the perturbative treatments of using the bare Green’s function, known as the Thouless criterion for superconductivity Thouless60 ; NSR85 and the Stoner criterion for ferromagnetism Stoner38 ; Blundell01 , for example. they are also identical with the mean-field estimations of critical temperatures as well as those based on random-phase approximations IKK63 ; BS66 ; DE66 . When making Green’s function self-consistent and incorporating terms other than the logarithmic contributions, however, each singular point generally becomes different from the corresponding transition point with a broken symmetry, as shown below both analytically and numerically. It also has not been clarified what happens after the singular points are reached and how to describe the phases theoretically. These are the issues to be studied here.
This paper is organized as follows. In Sect. 2, we elaborate the hidden singularities and introduce a procedure to handle them theoretically. In Sect. 3, we present numerical results for the negative-$U$ Hubbard model. In Sect. 4, we provide concluding remarks. We use the units of $\hbar=k_{\rm B}=1$.
@@ II.1 Luttinger-Ward functional
One of the theoretical frameworks suitable for the present purposes is the Luttinger-Ward expression of the grand thermodynamic potential $\Omega$ given as a functional of Green’s function $\hat{G}$ LW60 ; DDM64 ; DDM64-2 ; HRCZ07 ; Kita11 . The stationarity condition of $\Omega$ in terms of $\hat{G}$ yields a self-consistent equation for $\hat{G}$, which has been known, for example, as the Dyson, the Schwinger-Dyson, or Dyson-Gor’kov equation. Being nonlinear in $\hat{G}$, it can describe ordered phases with spontaneous broken symmetries continuously from the normal state on the same footing. Indeed, the ordered phases are characterized by the emergence of some novel component in the matrix $\hat{G}$ with particle-hole and spin degrees of freedom, such as the anomalous Green’s function for superconductivity and spin polarization in $\hat{G}$ for ferromagnetism. This Luttinger-Ward functional with full particle-hole and spin degrees of freedom can be written as Kita11
@@
where $\beta$ is the inverse temperature, $\hat{G}_{0}$ is the noninteracting Green’s function, $\hat{\Sigma}$ is the self-energy, and Tr denotes the trace over the arguments of $\hat{G}$. The stationarity condition of Eq. (1) with respect to $\hat{G}$ yields the Schwinger-Dyson equation:
@@
These are formally the exact expressions, and practical calculations can be performed on the basis of some approximate expression for $\Phi\!=\!\Phi[\hat{G}]$, obeying Eq. (2) similarly to the exact theory. The scheme of Eq. (2) has also been known as conserving approximations of Kadanoff and Baym KB62 or $\Phi$-derivable approximations Baym62 . The potential $\Omega$ as a functional of $\hat{G}$ is alternatively called the effective action in relativistic quantum field theory CJT74 ; Weinberg95 .
@@ II.2 Normal state
To elucidate the hidden singularities that lie near phase transitions with broken symmetries, we specifically consider the negative-$U$ Hubbard model at a low electron density $\bar{n}\!=\!N/V$, which has negligible lattice anisotropy and definitely exhibit $s$-wave superconductivity at low temperatures. We study this system on the basis of the fluctuation-exchange approximation for superconductivity (FLEX-S approximation), whose $\Phi$ is given diagrammatically in Fig. 1 Kita11 . We first focus on the normal state, where the series reduces to the $\Phi$ functional of the standard fluctuation-exchange (FLEX) approximation:BSW89 ; BS89 ; Kita11
@@
Here, $\vec{q}\equiv({\bm{q}},i\omega_{\ell})$ indicates the four-momentum consisting of the wavevector ${\bm{q}}$ and boson Matsubara frequency $\omega_{\ell}$. The functions $\chi_{GG}(\vec{q}\,)$ and $\chi_{G\bar{G}}(\vec{q}\,)$ are defined in terms of the normal-state Green’s function $G(\vec{k})$ and its time-reversed partner $\bar{G}(\vec{k})\!\equiv\!G(-\vec{k})$ generally by
@@
where $\vec{k}\!\equiv\!({\bm{k}},i\varepsilon_{n})$ with $\varepsilon_{n}$ denoting the fermion Matsubara frequency; this function satisfies
@@
as can be shown by a change in the summation variable $\vec{k}$. The first term in Eq. (3) is the Hartree-Fock contribution, the second one is the particle-particle ladder series relevant to the superconducting fluctuations including the second-order contribution, and the third and fourth terms represent the spin and charge fluctuations, respectively, composed of the particle-hole ladder and ring diagrams BSW89 . The corresponding self-energy is obtained by the functional differentiation of Eq. (2); for the normal state that is symmetric in the spin and particle-hole spaces, the differentiation is simplified to $\Sigma_{\rm n}(\vec{k})\!=\!\frac{\beta}{2}\frac{\delta\Phi_{n}}{\delta G(\vec{k})}$ in terms of Eq. (3), yielding BSW89 ; BS89 ; Kita11 ; Yanase03 ; Kontani13
@@
Now, we are ready to discuss the hidden singularities in full detail. With $G$ replaced by $G_{0}$ in Eq. (5), the conditions of the vanishing denominators at $\vec{q}=0$ in the second and third terms reproduce the Thouless criterion $U\chi_{G_{0}\bar{G}_{0}}(\vec{0})\!=\!1$ for superconductivity NSR85 and the Stoner criterion $U\chi_{G_{0}G_{0}}(\vec{0})\!=\!1$ for ferromagnetism Blundell01 , respectively. However, the renormalized conditions $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ do not specify the phase transition points of respective broken symmetries. Indeed, by linearizing the self-consistent equation for the anomalous Green’s function $F$ in the FLEX-S approximation, we obtain the equation for the superconducting transition temperature $T_{\rm c}$ that the eigenvalue problem
@@
has the non-trivial anomalous self-energy $\Delta(\vec{k})$ due to the emergence of a zero eigenvalue, where $U^{\rm s}_{\vec{k}\vec{k}^{\prime}}$ is the pairing interaction with correlations given explicitly by Eq. (91) in Ref. Kita11, , i.e.,
@@
This condition for $T_{\rm c}$, which is also identical with the one that the superconducting correlation function diverges, is manifestly different from $U\chi_{G\bar{G}}(\vec{0})\!=\!1$. The same statement holds true in the case of ferromagnetism with $U\!>\!0$, as can be examined easily on the basis of the consideration of incorporating the spin polarization in Green’s function. Hence, we should allow for the possibility that $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and $U\chi_{GG}(\vec{0})\!=\!1$ are reached independently of the phase transitions with broken symmetries. The fact that $U\chi_{G_{0}\bar{G}_{0}}(\vec{0})\!=\!1$ and $U\chi_{G_{0}G_{0}}(\vec{0})\!=\!1$ specify the mean-field $T_{\rm c}$ suggests that the renormalized ones are realized near the phase transition points with broken symmetries.
@@
$\frac{1}{\beta}\sum_{\bm{q}}\left\{\ln\left[1-U\chi_{G\bar{G}}(q,0)\right]+U\chi_{G\bar{G}}(q,0)\right\},$ (8)
where $U\chi_{G\bar{G}}(q,0)$ is real, as can be shown from Eq. (4). Since Eq. (8) is part of the grand thermodynamic potential that takes a real value, the inequality $11-U\chi_{G\bar{G}}(q,0)\!\geq\!$ 1 should hold. In other words, the series expansion of Eq. (8) does not converge for $U\chi_{G\bar{G}}(q,0)>1$. Thus, we are led to the conclusion that $U\chi_{G\bar{G}}(\vec{0})$ and $U\chi_{GG}(\vec{0})$ must have the common bound 1, at which they should be pinned after reaching it. The statement holds true not only for $q\!=\!0$ of isotropic systems but also for any value of ${\bm{q}}$ even in anisotropic systems; we call $U\chi_{G\bar{G}}({\bm{q}},0)\!=\!1$ and $U\chi_{GG}({\bm{q}},0)\!=\!1$ hidden singularities. It is shown in AppendixA that the hidden singularities can be defined generally beyond the FLEX-S approximation.
@@
We propose a resolution to this issue within the variational framework of Eqs. (1) and (2) by using the method of the Lagrange multiplier so as not to affect $\Omega$ explicitly. Specifically for the negative-$U$ case where $U\chi_{G\bar{G}}(\vec{0})$ is relevant, we add to the $\Phi$ functional the following term that vanishes for $U\chi_{G\bar{G}}(\vec{0})=1$:
$\varDelta\Phi_{\lambda}^{\rm n}\equiv\lambda_{\rm n}N\bigl{[}1-U\chi_{G\bar{G}}(\vec{0})\bigr{]},$ (9)
@@
where the number of particles, $N$, is incorporated in the prefactor to make the Lagrange multiplier $\lambda_{\rm n}$ intensive; note that $\lambda_{\rm n}$ has the dimension of energy. Performing the differentiation of Eq. (2) with Eq. (9) added to $\Phi$, we obtain the expression of the self-energy for the pinned phase of $U\chi_{G\bar{G}}(\vec{0})=1$ in terms of Eq. (5) as
$\Sigma_{\lambda}^{\rm n}(\vec{k})=\Sigma_{\rm n}(\vec{k})+\bar{n}U\lambda_{\rm n}G(-\vec{k})$ (10)
@@
with $\bar{n}\!\equiv\!N/V$. The additional second term stemming from Eq. (9) is characteristic of the pinned phase, which corresponds to the condensate fraction of the ideal Bose-Einstein condensation Leggett06 . It has an unusual structure called one-particle reducible (1PR), i.e., the structure that becomes disconnected upon cutting a line,CJT74 which originates from the two-particle reducible structure of Eq. (9). The parameter $\lambda_{\rm n}$ in Eq. (9) enables us to control the value of $U\chi_{G\bar{G}}(\vec{0})$ and should be chosen in such a way that $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ is met by self-consistently solving the Dyson equation:
$G(\vec{k})\!=\!\bigl{[}G_{0}^{-1}(\vec{k})\!-\!\Sigma_{\lambda}^{\rm n}(\vec{k})\bigr{]}^{-1}.$ (11)
This completes the formulation for the normal state.
@@
The parameter $\lambda_{\rm n}$ is finite only for the pinned phase of $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ and is expected to grow almost linearly as a function of temperature initially. This linear change in the structure of the self-energy need not be considered in performing the first derivatives of $\Omega$ because of stationarity condition (2). However, it does bring singularities in the second derivatives of $\Omega$ such as specific heat, similarly to those caused by the emergences of $\Delta(\vec{k})$ for superconductivity and the spin-polarized self-energy for ferromagnetism near $T_{\rm c}$. Thus, the point $U\chi_{G\bar{G}}(\vec{0})\!=\!1$ may be regarded as specifying a kind of transition without apparent broken symmetries. In this respect, note that Green’s function with the self-energy of Eq. (10) cannot be approximated by the quasiparticle form of the Fermi-liquid theory,AGD63 $G(\vec{k})\approx a/(i\varepsilon_{n}-\xi_{\bm{k}})$; hence, the singular point may be identified as a non-Fermi-liquid transition. Detailed studies on observable singularities remain to be performed in the future.
Several comments are in order on the hidden singularities and also on Eq. (9) introduced here to treat them. There are two possibilities for the hidden singularities. The first possibility is that the singularities specify the limits of the Luttinger-Ward formalism beyond which it cannot be used, with no real physical phenomena connected to them; the statement implies that there is a large domain, especially in and near ordered phases with broken symmetries, where the self-consistent perturbative scheme is ineffective. The second possibility is that the singularities exist physically, as assumed in the bare perturbation scheme with the names of Thouless and Stoner criteria, which should be handled appropriately once they have been reached, either by extending the Luttinger-Ward formalism itself or by relying on some alternative formalism suitable for strongly correlated systems; since the hidden singularities are generally distinct from the second-order transition points with broken symmetries, critical fluctuations should be irrelevant to resolving the issue. We have taken the second possibility and developed the method given by Eq. (9) to take care of the singularities within the Luttinger-Ward formalism, through our numerical calculations for the negative-$U$ Hubbard model elaborated below in Sect. 3. This method yields a definite theoretical prediction, which can be tested by experiments, that the reaching a singularity can be detected as singularities in the second derivatives of the thermodynamic potential such as specific heat, as mentioned above. Moreover, the resulting Eq. (10) with $G(-\vec{k})$ naturally explains the emergence of a pole in the self-energy, which has been discussed in terms of the pseudogap behaviors in high-$T_{\rm c}$ cuprates phenomenologicallyNorman07 and also in the negative-$U$ Hubbard model found numerically by a completely different approach of the dynamical mean-field theory.PB15 ; Sakai15 These are some of the plausible features of our approach with Eq. (9).